Solve the following set of equation by both graphing method and also by addition.
Prepare a paragraph comparing and contrasting the two methods.

3x+4y = -5
5x+6y = -7
Please :)

Answer by stanbon(75887)

Solve the following set of equation by both graphing method and also by addition.
Prepare a paragraph comparing and contrasting the two methods.
3x+4y = -5
5x+6y = -7
----------------------------
Solve by elimination:
Multiply thru 1st equation by 5
Multiply thru 2nd equation by 3
-------
15x + 18y = -21
15x + 20y = -25
---
Subtract from 1st from 2nd and sol solve for "y":
y = -4
---
Substitute into 3x+4y=-6 to solve for "x":
3x +4(-4) = -6
3x = 10
x = 2
====================================================
3x+4y = -5
5x+6y = -7
------
Solve by graphing:
y = (-3/4)x -(5/4)
Let x = 0, then y = -5/4 giving you point (0,-5/4)
Let x = 4, then y = -17/4 giving you point (4,-17/4)
Plot the points and draw a line thru them.
----------------------
Solve by graphing:
y = (-5/6)x -(7/6)
Let x = 0, then y = -7/6 giving you point (0,-17/6)
Let x = 6, then y = -37/6 giving you point (6,-36/6)
Plot the points and draw a line thru them.
-------------------------
Graph and find the point is intersection; it should be (2,-4)
===========================

=============================
Cheers,
Stan H.